Two people start driving toward each other on the same highway at the same time from towns 420 miles apart. one car travels 15 mph faster then the other. what is the rate of each car if they meet in 4 hours?

answer: the first car travels 60 mph; the second car travels at 45 mph.

need a full break down of how my book got this answer i know that i have to use r=d/t

Answer by Mathtut(3670)

d=r*t
:
d=(r+15)4...eq 1---->d=4r+60.......fast car
420-d=r(4)..eq 2---->d=-4r+420.....slow car
:
4r+60=-4r+420
:
8r=360
:
r=45 slow car
:
r+15=60...fast car
: